# Optical probes of two-component pairing states in transition metal dichalcogenides

### v1

@@ Abstract
Certain transition metal dichalcogenide heterostructures based on 2H-NbSe2 and 4Hb-TaS2 have recently displayed surprising signatures of unconventional superconductivity. While the pairing channel remains unknown, it has been argued that spin fluctuations should lead to a two-component $E^{\prime}$ pairing state which is compatible with some experimental features. Exploiting the particular multi-orbital character of the Fermi surface and the presence of Ising spin-orbit coupling which enable finite optical conductivity in the clean limit, in this work we predict clear-cut optical signatures of the chiral and nematic ground states of the $E^{\prime}$ pairing. In particular, we show how the spontaneous breaking of time-reversal and threefold symmetry are reflected in the optical Hall and anisotropic conductivities, respectively, while different spectral features can be connected with the momentum dependence of the gap functions. Our work provides a fingerprint that can be measured experimentally, constraining the pairing channel in TMD superconductors, and will help determine whether superconductivity is topological in these systems.
@@
Introduction–– Metallic transition metal dichalcogenides (TMDs) MX2 (M=Nb,Ta, X=S, Se) Wilson et al. (1975), believed to be standard BCS superconductors in their bulk 2H polytype, have recently shown signatures of unconventional superconductivity Norman (2011) when made into thin films and heterostructures, calling for a better understanding of their pairing symmetries and mechanisms. For example, NbSe2 thin films display spontaneous breaking of threefold symmetry in the superconducting state under in-plane magnetic fields (indicative of a two-component order parameter) Hamill et al. (2021); Cho et al. (2022), as well as superconducting collective modes Wan et al. (2022) and anomalous resilience to magnetic fields Kuzmanović et al. (2022). Topological superconductivity, necessarily unconventional, has also been reported in monolayer CrBr3 on bulk NbSe2 Kezilebieke et al. (2020). 4Hb-TaS2 Di Salvo et al. (1973), which can be seen as a heterostructure of H and T polytypes, shows spontaneous breaking of time-reversal symmetry (TRS) at the critical superconducting temperature $T_{c}$ Ribak et al. (2020), spontaneous vortices in the superconducting state Persky et al. (2022), superconducting edge modes Nayak et al. (2021) and transport evidence of a two-component order parameter, which also includes spontaneous breaking of threefold symmetry Almoalem et al. (2022); Silber et al. (2024). A chiral superconducting state has also been reported in 2H-TaS2 intercalated with chiral molecules Wan et al. (2023).
@@
The spontaneous breaking of threefold or time-reversal symmetries suggest a two-component SC order parameter, which may condense into nematic or chiral states depending on the experiment, possibly competing with a conventional pairing channel. The microscopic origin of such two-component pairing remains unknown, but it is strongly constrained by the presence of Ising spin-orbit coupling (SOC) Xi et al. (2016); Sohn et al. (2018); De la Barrera et al. (2018); Wickramaratne et al. (2020), induced by a crystal structure which breaks inversion symmetry but preserves a horizontal mirror $\sigma_{h}$. The pairing problem for H-TMDs has been studied at length He et al. (2018); Möckli and Khodas (2018); Shaffer et al. (2020); Margalit et al. (2021); Dentelski et al. (2021); Hörhold et al. (2023); Liu et al. (2024), in particular in the presence of two-component order parameters Hsu et al. (2017); Chen et al. (2019); Lane and Zhu (2022), but without identifying a microscopic interaction that makes these channels attractive. Previous works had however identified that monolayer H-TMDs are near a magnetic instability towards an antiferromagnetic state with in-plane polarization Zheng et al. (2018); Wickramaratne et al. (2020); Das and Mazin (2021); Costa et al. (2022), which leads to strong transverse paramagnons that can mediate unconventional superconductivity. A more recent calculation reveals that such fluctuations favor an $E^{\prime}$ triplet state in the presence of Ising SOC Roy et al. (2024); Siegl et al. (2024), although it remains to be established whether this solution is robust in the presence of additional electron-phonon coupling Das et al. (2023). This $E^{\prime}$ channel, which pairs electrons of opposite spins (and is therefore unaffected by Ising SOC), is a two-component representation of the point group that can condense in nematic $(p_{x},p_{y})$ or chiral $p_{x}+ip_{y}$ ground states. It is thus a strong candidate to explain the phenomenology observed in TMD systems, and an experimental probe that can distinguish it from other channels would be desirable.
@@
It has recently been realized that the optical conductivity is a useful probe of the superconducting state of multiband (or multiorbital) systems Ahn and Nagaosa (2021a). While in single-band superconductors the optical conductivity is essentially disorder-driven and vanishes in the clean limit, clean multiband superconductors have finite optical conductivity, which is sensitive to the pairing channel and is connected to the superfluid weight Ahn and Nagaosa (2021b); Chen and Huang (2021). It is also ideally suited to probe the difference between nematic and chiral ground states. Conductivity anisotropy $\sigma_{xx}-\sigma_{yy}$ is sensitive to the breaking of $C_{3}$ symmetry Xu et al. (2022), while optical Hall conductivity is only allowed in chiral systems which break time-reversal symmetry Taylor and Kallin (2012); Wang et al. (2014, 2017). Previous work has considered the non-linear optical response of non-centrosymmetric superconductors Xu et al. (2019); Tanaka et al. (2023), and recently both linear and non-linear conductivities were computed for the H-TMD three band model Raj et al. (2024) with the single-component pairing states proposed in Ref. [Margalit et al., 2021]. The $E^{\prime}$ state, currently a strong candidate for unconventional pairing Roy et al. (2024); Siegl et al. (2024), was not considered.
@@
Model and pairing channels–– The H polytype of TMDs MX2 (M=Nb,Ta, X=S, Se) crystallizes in a trigonal prismatic crystal structure (Fig 1 (a)) with $D_{3h}$ point-group symmetry. The Fermi surface derives from the ${d_{z^{2}},d_{x^{2}-y^{2}},d_{xy}}$ orbitals of the transition metalMöckli and Khodas (2018). We consider a spinful three-band tight-binding model with up to third-nearest neighbours hoppingsLiu et al. (2013) in the orbital basis $\{d_{z^{2}},d_{x^{2}-y^{2}},d_{xy}\}$ which reads
@@
where $h^{\mathrm{TNN}}_{\bm{k}}$ is the tight-binding Hamiltonian with up to third nearest-neighbours hoppings, $\lambda$ is the SOC strength, and $L_{z}$ is the matrix of the z-component of the angular momentum in the orbital basis $\{d_{z^{2}},d_{x^{2}-y^{2}},d_{xy}\}$ (See Ref. [Liu et al., 2013] for details and Fig. 1 (b) for the resulting band structure.).
@@
The different pairing channels are classified as irreducible representations (irreps) of the $D_{3h}=M_{z}\otimes C_{3v}$ point group, generated by mirror planes perpendicular to the $x$-axis $M_{x}$ and $z$-axis $M_{z}$, and threefold rotations $C_{3}$ around the $z$-axis. This group contains the $M_{z}$-even irreps $A_{1}^{\prime}$, $A_{2}^{\prime}$, $E^{\prime}$ and $M_{z}$-odd irreps $A_{1}^{\prime\prime}$, $A_{2}^{\prime\prime}$, $E^{\prime\prime}$. Since the orbitals that contribute to the bands at the Fermi level are all $M_{z}$-even and we only consider in-plane pairing, odd $M_{z}$ parity can only come from the spin part of the pairing. Singlet pairing $\Delta=i\sigma_{y}$ and out-of-plane triplet $\Delta=i\sigma_{y}\sigma_{z}=\sigma_{x}$ are $M_{z}$-even pairings (which correspond to opposite spin pairing), while in-plane triplets $\Delta=i\sigma_{y}(\sigma_{x},\sigma_{y})$ are $M_{z}$-odd pairings (which correspond to equal-spin pairing). Since $D_{3h}$ does not contain the inversion operator, singlet and triplet channels can generally mix due to SOC, and channels are only classified by the global (spin and orbital) symmetry of the pairing.
@@
In addition to being unaffected by Ising SOC and favored by spin fluctuations, the $E^{\prime}$ channel has a number of appealing features. First, as a two-component order parameter it admits both chiral and nematic ground states, in line with the phenomenology in 4Hb-TaS2. In addition, this state has interesting topological features. $M_{z}$-even pairing with Ising SOC implies spin rotation invariance around $z$ direction, which puts this system in topological class AIII Schnyder et al. (2008), featuring $Z$ invariants in 1D that can protect nodal points in the case of nematic ground states. For chiral ground states, each mirror sector realizes a Chern number $C=3$ superconductor. In the case of triplet $E^{\prime}$ chiral state, a $C=6$ superconductor is realized, which could also be consistent with the edge state phenomenology in 4Hb-TaS2Nayak et al. (2021) and CrBr3/NbSe2Kezilebieke et al. (2020).
Figure 2: Band structure of the BdG Hamiltonian in Eq.(6) for 4Hb-TaS2 with the superconducting pairings (a) $\Delta_{p_{x}}$, (b) $\Delta_{p_{y}}$, and (c) $\Delta_{p+ip}$ for $\Delta=0.1$ eV. The dark (light) gray shadowing corresponds to the minimal (maximal) size of the gap in the band structure along the high-symmetry directions. The insets correspond to the gap size across the Brillouin Zone for each pairing, where lightest colors are closer to zero gap, and indicate the high-symmetry directions corresponding to the band structure plots. The dotted-dashed line in (c) corresponds to the energy scale associated with the Van Hove singularity giving rise to the secondary peak in $\sigma_{xx}$ and $\sigma^{H}_{xy}$.
@@
BdG quasiparticle band structure –– For concreteness we consider a single layer of H-TaS2 and use the normal state tight-binding parameters reported in Ref. [Margalit et al., 2021]. Since to resolve band structure and optical features with realistic values of the superconducting gap ($\sim$1 meV) would require a very fine k-mesh with high computational cost, for illustrative purposes we consider pairings of an artificially large strength $\Delta=0.1$ eV, keeping in mind that experimental features are to be expected at the real values of the superconducting gap. The three different ground states with $E^{\prime}$ pairing $\Delta_{p_{x},p_{y},p+ip}$ lead to the three different band structures in Fig. 2. The $\Delta_{p_{x}}$ pairing is gapless along the $\Gamma M^{\prime}$ ($k_{x}=0$) direction, while the $\Delta_{p_{y}}$ pairing is gapless along the $\Gamma K$ ($k_{y}=0$) direction. In both cases we see that $\Gamma M$ and $\Gamma M^{\prime}$ are not equivalent, a consequence of the breaking of $C_{3}$ symmetry. For both nematic phases, the maximum gap is approximately $1.6\Delta$ along the $\Gamma K$ ($\Gamma M^{\prime}$) directions for $\Delta_{p_{x}}$ ($\Delta_{p_{y}}$). The $\Delta_{p+ip}$ phase, on the other hand, restores $C_{3}$ symmetry (combined with a gauge transformation) so the energies are $C_{3}$-symmetric, and it exhibits a full gap ranging in size from $0.43\Delta$ along the $MK$ direction to $1.6\Delta$ along the $\Gamma M$ direction.
Optical conductivity–– Since we consider exclusively $k-$independent pairings, Peierls substitution couples the electromagnetic field $\vec{k}\to\vec{k}-e\vec{A}$ through the normal-state Hamiltonian $h_{k}$, and the conductivity can be computed using standard Kubo formalismAhn and Nagaosa (2021a):

### v2

@@ Abstract
Signatures of unconventional superconductivity have been recently observed in certain transition metal dichalcogenides (TMDs), including 4Hb-TaS2 and monolayer 2H-NbSe2. While the pairing channel remains unknown, it has been argued that spin fluctuations can stabilize pairing in the two-component $E^{\prime}$ channel, a $p$-wave spin-triplet state which could be consistent with some of the reported signatures. Exploiting the particular multi-orbital character of the Fermi surface and the presence of Ising spin-orbit coupling, which enable finite optical conductivity in the clean limit, in this work we predict clear-cut optical signatures to detect and distinguish the chiral and nematic ground states of the $E^{\prime}$ pairing. We quantify how nematic $E^{\prime}$ states produce a diagonal anisotropy $\sigma_{xx}\!\neq\!\sigma_{yy}$ due to the broken threefold symmetry ($C_{3}$), while chiral $E^{\prime}$ states yield a finite optical Hall conductivity $\sigma_{xy}^{H}$ due to broken time-reversal symmetry, and find both signals could be detected in current experiments. For instance, for realistic gaps in the meV range, we predict a relative anisotropy $\Delta\sigma/\sigma\sim 10^{-5}$ in the nematic states, and a polar Kerr rotation of $\theta_{K}\!\sim\!10^{-5}$ rad in the chiral states. These symmetry fingerprints provide a practical route to distinguish nematic and chiral superconducting order in TMD superconductors.
I Introduction
@@
Metallic transition metal dichalcogenides (TMDs) MX2 (M=Nb,Ta, X=S, Se) Wilson et al. (1975), believed to be standard BCS superconductors in their bulk 2H polytype, have recently shown signatures of unconventional superconductivity Norman (2011) when made into thin films and heterostructures, calling for a better understanding of their pairing symmetries and mechanisms. For example, NbSe2 thin films display spontaneous breaking of threefold symmetry in the superconducting state under in-plane magnetic fields  Hamill et al. (2021); Cho et al. (2022), as well as superconducting collective modes Wan et al. (2022) and anomalous resilience to magnetic fields Kuzmanović et al. (2022). Evidence for topological edge states, which originate from unconventional pairing, has also been reported in monolayer CrBr3 on bulk NbSe2 Kezilebieke et al. (2020). 4Hb-TaS2 Di Salvo et al. (1973), which can be seen as a heterostructure of H and T polytypes, shows anomalous muon spin rotation at the critical temperature $T_{c}$ which has been claimed to originate from time-reversal symmetry (TRS) breakingRibak et al. (2020). Other experiments reported spontaneous vortices in the superconducting state Persky et al. (2022), superconducting edge modes Nayak et al. (2021) and anomalous Little-Parks oscillations and upper critical field anisotropy Almoalem et al. (2024); Silber et al. (2024). A chiral superconducting state has also been reported in 2H-TaS2 intercalated with chiral molecules Wan et al. (2024). Since a two-component order parameter may condense into nematic or chiral superconducting states, thus breaking threefold and time-reversal symmetries, respectively, these experiments represent tantalizing indirect evidence of two-component pairing. Demonstrating its existence as a ground state stable phase, however, remains a major challenge, because so far the evidence of broken TRS is only indirect and broken threefold symmetry only occurs in the presence of an applied field.
@@
The potential candidate channels for unconventional pairing in these TMD systems are strongly constrained by the presence of Ising spin-orbit coupling (SOC) Xi et al. (2016); Sohn et al. (2018); De la Barrera et al. (2018); Wickramaratne et al. (2020), induced by a crystal structure which breaks inversion symmetry but preserves a horizontal mirror $\sigma_{h}$. The pairing problem for H-TMDs has been studied at length He et al. (2018); Möckli and Khodas (2018); Shaffer et al. (2020); Margalit et al. (2021); Dentelski et al. (2021); Hörhold et al. (2023); Liu et al. (2024), in particular in the presence of two-component order parameters Hsu et al. (2017); Chen et al. (2019); Lane and Zhu (2022), but without identifying a microscopic interaction that makes these channels attractive. Previous works had however identified that monolayer H-TMDs are near a magnetic instability towards an antiferromagnetic state with in-plane polarization Zheng et al. (2018); Wickramaratne et al. (2020); Das and Mazin (2021); Costa et al. (2022), which leads to strong transverse paramagnons that can mediate unconventional superconductivity. A more recent calculation reveals that such fluctuations favor the two-component $E^{\prime}$ triplet state in the presence of Ising SOC Roy et al. (2024); Siegl et al. (2025), although it remains to be established whether this solution is robust in the presence of additional electron-phonon coupling Das et al. (2023); Roy et al. (2025). This $E^{\prime}$ channel, which pairs electrons of opposite spins, is a $p$-wave triplet unaffected by Ising SOC and can condense in nematic $(p_{x},p_{y})$ or chiral $p_{x}+ip_{y}$ ground states. It is thus a strong candidate to explain the phenomenology observed in TMD systems.
Since the evidence of TRS and threefold symmetry breaking in these TMDs is so far indirect, an experimental probe that is directly sensitive to both symmetries in the superconducting state would be desirable to establish the existence of the $E^{\prime}$ ground state. In this work, we choose the linear optical conductivity as such a probe. Indeed, it is well known that conductivity anisotropy $\sigma_{xx}\!\neq\!\sigma_{yy}$, which leads to observable features like linear dichroism and birrefringence, becomes enabled when perpendicular rotation axes $C_{n}$ with $n>2$ are absent Rodger and Nordén (1997); Newnham (2004); Mirri et al. (2016); Xu et al. (2022), like in nematic states. In addition, Onsager relations dictate that the antisymmetric (or Hall) part of the conductivity, which leads to magnetic circular dichroism and polar Kerr effect Argyres (1955); Xia et al. (2006); Kapitulnik et al. (2009); Schemm et al. (2014); Taylor and Kallin (2012); Wang et al. (2014, 2017), is only allowed when time-reversal and vertical mirror symmetries are broken, like in the chiral superconducting state. Optical conductivity is therefore ideally suited to probe the $E^{\prime}$ state.
@@
Historically, the optical conductivity of superconductors has been understood in the Mattis–Bardeen, single-band dirty limit, which is essentially disorder-driven and leads to vanishing conductivity in the absence of scattering. But more recently it has been realized that clean multiband superconductors have finite optical conductivity Ahn and Nagaosa (2021a), which is sensitive to the pairing channel and is connected to the superfluid weight Ahn and Nagaosa (2021b); Chen and Huang (2021). Previous work has considered the non-linear optical response of non-centrosymmetric superconductors Xu et al. (2019); Tanaka et al. (2023), and recently both linear and non-linear conductivities were computed for the H-TMD three band model Raj et al. (2024) with the single-component pairing states proposed in Ref. [Margalit et al., 2021]. The $E^{\prime}$ state Roy et al. (2024); Siegl et al. (2025) was not considered.
@@
II Model and pairing channels
The H polytype of monolayer TMDs MX2 (M=Nb,Ta, X=S, Se) crystallizes in a trigonal prismatic crystal structure (Fig 1 (a)) with $D_{3h}$ point-group symmetry. The Fermi surface derives from $d_{z^{2}},d_{x^{2}-y^{2}},d_{xy}$}-like orbitals centered at the transition metal positionsMöckli and Khodas (2018). We consider a spinful three-band tight-binding model with up to third-nearest neighbours hoppingsLiu et al. (2013) in the orbital basis $\{d_{z^{2}},d_{x^{2}-y^{2}},d_{xy}\}$ which reads
@@ II Model and pairing channels
where $h^{\mathrm{TNN}}_{\bm{k}}$ is the tight-binding Hamiltonian with up to third nearest-neighbours hoppings, $\lambda$ is the SOC strength, and $L_{z}$ is the matrix of the z-component of the angular momentum in the orbital basis $\{d_{z^{2}},d_{x^{2}-y^{2}},d_{xy}\}$. See Ref. [Liu et al., 2013] for parameter details and Fig. 1(b) for the resulting band structure.
@@
The different pairing channels are classified as irreducible representations (irreps) of the $D_{3h}=M_{z}\otimes C_{3v}$ point group, generated by mirror planes perpendicular to the $x$-axis $M_{x}$ and $z$-axis $M_{z}$, and threefold rotations $C_{3}$ around the $z$-axis. This group contains the $M_{z}$-even irreps $A_{1}^{\prime}$, $A_{2}^{\prime}$, $E^{\prime}$ and $M_{z}$-odd irreps $A_{1}^{\prime\prime}$, $A_{2}^{\prime\prime}$, $E^{\prime\prime}$. Since the orbitals that contribute to the bands at the Fermi level are all $M_{z}$-even and we only consider intralayer pairing, odd $M_{z}$ parity can only come from the spin part of the pairing. Singlet pairing $\Delta=i\sigma_{y}$ and out-of-plane triplet $\Delta=i\sigma_{y}\sigma_{z}=\sigma_{x}$ are $M_{z}$-even pairings (which correspond to opposite spin pairing), while in-plane triplets $\Delta=i\sigma_{y}(\sigma_{x},\sigma_{y})$ are $M_{z}$-odd pairings (which correspond to equal-spin pairing). Since $D_{3h}$ does not contain the inversion operator, singlet and triplet channels can generally mix due to SOC, and channels are only classified by the global (spin and orbital) symmetry of the pairing.
@@
In addition to being unaffected by Ising SOC and favored by spin fluctuations, the $E^{\prime}$ channel has a number of appealing features. First, as a two-component order parameter it admits both chiral and nematic ground states, in line with the phenomenology in 4Hb-TaS2Ribak et al. (2020); Persky et al. (2022); Nayak et al. (2021); Almoalem et al. (2024); Silber et al. (2024). In addition, this state has interesting topological features. $M_{z}$-even pairing with Ising SOC implies spin rotation invariance around $z$ direction, which puts this system in topological class AIII Schnyder et al. (2008), featuring $Z$ invariants in 1D that can protect nodal points in the case of nematic ground states. For chiral ground states, each mirror sector realizes a Chern number $C=3$ superconductor. In the case of triplet $E^{\prime}$ chiral state, a $C=6$ superconductor is realized, which could also be consistent with the edge state phenomenology in 4Hb-TaS2Nayak et al. (2021) and CrBr3/NbSe2Kezilebieke et al. (2020).
Figure 2: Band structure of the BdG Hamiltonian in Eq. (6) for 4Hb-TaS2 with the superconducting pairings (a) $\Delta_{p_{x}}$, (b) $\Delta_{p_{y}}$, and (c) $\Delta_{p+ip}$ for $\Delta=0.1$ eV. The dark (light) gray shading marks the minimal (maximal) gap along high-symmetry directions. Insets show the gap magnitude across the Brillouin zone; light colors indicate smaller gaps and insets mark the paths used in the main panels. The dotted-dashed line in (c) marks the energy associated with the van Hove singularity that produces the secondary peak in the absorptive components $\mathrm{Re}\sigma_{xx}$ and $\mathrm{Im}\sigma^{H}_{xy}$.
@@ III BdG quasiparticle band structure
For concreteness, we consider a single layer of H-TaS2 and use the normal state tight-binding parameters reported in Ref. [Margalit et al., 2021]. Band structure effects and optical features of the superconducting state with realistic values of the superconducting gap ($\sim$1 meV) are, in principle, quite small. For illustrative purposes we consider pairings of an artificially large strength $\Delta=0.1$ eV, keeping in mind that experimental features are to be expected at the real values of the superconducting gap. We will focus on realistic values of $\Delta$ and the scaling of the relevant features later on. The three different ground states with $E^{\prime}$ pairing $\Delta_{p_{x},p_{y},p+ip}$ lead to the three different band structures in Fig. 2. The $\Delta_{p_{x}}$ pairing is gapless along the $\Gamma M^{\prime}$ ($k_{x}=0$) direction, while the $\Delta_{p_{y}}$ pairing is gapless along the $\Gamma K$ ($k_{y}=0$) direction. In both cases we see that $\Gamma M$ and $\Gamma M^{\prime}$ are not equivalent, a consequence of the breaking of $C_{3}$ symmetry. For both nematic phases, the maximum gap is approximately $1.6\Delta$ along the $\Gamma K$ ($\Gamma M^{\prime}$) directions for $\Delta_{p_{x}}$ ($\Delta_{p_{y}}$). The $\Delta_{p+ip}$ phase, on the other hand, restores $C_{3}$ symmetry (combined with a gauge transformation) so the energies are $C_{3}$-symmetric, and it exhibits a full gap ranging in size from $0.43\Delta$ along the $MK$ direction to $1.6\Delta$ along the $\Gamma M$ direction.
IV Optical conductivity
To compute the optical conductivity, defined from $J_{i}=\sigma_{ij}E_{j}$, we employ the standard Kubo formalism as used for superconductors in Ref. [Ahn and Nagaosa, 2021a]. Since we consider $k$-independent pairings, Peierls substitution couples the electromagnetic field $\vec{k}\to\vec{k}-e\vec{A}$ through the normal-state Hamiltonian $h_{k}$. The conductivity tensor $\sigma_{ij}$ can be separated into symmetric and antisymmetric (Hall) parts. For the symmetric part we only consider the diagonal components. In the main text we report the absorptive part of the conductivity $\sigma_{ij}^{\rm abs}=\tfrac{1}{2}(\sigma_{ij}+\sigma_{ji}^{*})$. For the symmetric part, the absorptive part is real $\sigma_{ii}^{\rm abs}={\rm Re}\sigma_{ii}$, while for the Hall part $\sigma^{H}_{xy}=\frac{1}{2}\left(\sigma_{xy}-\sigma_{yx}\right)$, the absorptive part is imaginary $\sigma_{ij}^{H,\rm abs}={\rm Im}\sigma_{ij}^{H}$. The reactive parts $\sigma_{ij}^{\rm rea}=\tfrac{1}{2}(\sigma_{ij}-\sigma_{ji}^{*})$ can be computed through Kramers-Kronig relations (see Appendices A and  B). The absorptive, diagonal part is explicitly given by
${\rm Re}\sigma_{ii}=\frac{\pi e^{2}}{2\omega V}\sum_{m\neq n}v^{i}_{nm}v^{i}_{mn}f_{nm}\delta(\varepsilon_{n}-\varepsilon_{m}+\omega),$ (8)
while the absorptive Hall part is given by